OpenStudy (anonymous):

Two electrically-charges spheres are suspended from insulated threads a certain distance from each other. There is a certain amount of electrostatic force between them. Describe specifically (not just increase or decrease) what happens to this force in each of the scenarios below (1 pt. ea.): The charge on one sphere is reduced by half The charge on both spheres is doubled The distance between the spheres is increased by a factor of three The distance between the sphere is decreased to one-fourth The charge of each sphere is doubled and the distance between them is doubled

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

let's suppose to have 2 charges, namely Q_1 and Q_2, separated by a distance d. Then the interaction force F (its magnitude) is equal to: \[F = k\frac{{{Q_1}{Q_2}}}{{{d^2}}}\]

OpenStudy (michele_laino):

now, we have: scenario #1 let's suppose that the new charge is Q_1/2, then the new force is: \[{F_1} = k\frac{{\left( {{Q_1}/2} \right){Q_2}}}{{{d^2}}} = k\frac{{{Q_1}{Q_2}}}{{2{d^2}}} = \frac{F}{2}\]

OpenStudy (michele_laino):

please do the same with other scenarios